OpenStudy (lemanual):

The sum of three consecutive integers is 267. What is the largest integer? PLEASE HELP ME

OpenStudy (anonymous):

Just divide 267 by 3. Then what do you get?

OpenStudy (anonymous):

so consecutive will be one bigger than x= x+1 and other consecutive will be one bigger than x+1=x+2 now sum of these 3 = 267 x+x+1+x+2=267 3x+3=267 3x=264 x=88 now we know that the biggest one is x+2 right? plus in the value of x 88+2=90 which is biggest

OpenStudy (anonymous):

Well, lets say the first number is n, so we get the equation n+(n+1)+(n+2)=267. This gives us 3n+3=267, or 3n=264. So n=264/3=88. Now we know the first number is 88, so the second is 89 and the third is 90.

OpenStudy (anonymous):

Then round it and you get 90.

OpenStudy (anonymous):

This way is much faster.

OpenStudy (lemanual):

Thanks so much guys can i ask you another one?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Sure

OpenStudy (lemanual):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax and x represents the original price of the item. : C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale? Remember, do not include the $ symbol.

OpenStudy (anonymous):

90

OpenStudy (lemanual):

Anyone knows this?